# Weakly Supervised Learning for Facial Behavior Analysis: A Review

### v3

@@ Abstract
Given the recent advances in deep learning (DL), and with sensor and computing technologies, there has been considerable progress in the development of systems that analyze facial behavior, evolving from systems that perform image analysis under controlled laboratory conditions, to those for video-based recognition under more challenging real-world conditions. However, some key challenges in real-world applications include a wide range of temporal dynamics of facial expressions, diverse capture conditions, and the limited amount of relevant data to train predictive models. DL models typically require training with large-scale datasets to provide a high level of performance, yet the collection and annotation of such datasets is a costly undertaking that relies on domain experts. Moreover, the annotation process is highly vulnerable to the ambiguity of expressions or AUs due to the bias induced by experts. This paper provides a comprehensive review of state-of-the-art weakly supervised learning (WSL) methods for facial behavior analysis based on images and videos. First, we introduce a taxonomy of relevant WSL scenarios, and then methods proposed for classification applications (using weak discrete/categorical labels), and for regression applications (using weak ordinal/dimensional labels). For both types of recognition applications (classification and regression), we provide a systematic review of state-of-the-art machine learning (ML) models for expression and AU recognition, and intensity estimation in each WSL scenario, along with their respective strengths and limitations. An overview of widely used experimental methodologies (public datasets and protocols) is also provided for the performance evaluation of these state-of-the-art models. Finally, our critical analysis of the methods, along with their experimental results, provides insight into the key challenges of WSL in different scenarios and applications and motivates new research directions to leverage weakly labeled data to address real-world facial behavior analysis problems. Our review indicates that WSL methods may provide a cost-effective approach to train robust ML/DL models for facial behavior analysis in real-world scenarios.
@@
Affective Computing, Weakly Supervised Learning, Facial Expression Recognition, Intensity Estimation, Action Unit Detection, Deep Learning.
1 Introduction
Facial Behavior Analysis is an emerging area of interest in computer vision and affective computing, with immense potential for many applications in human-computer interaction, sociable robots, autonomous-driving cars, etc. It was shown that only one-third of human communication is conveyed through verbal components and two-thirds of communication occurs through non-verbal components [1]. Although emotions can be conveyed through multiple modalities, facial behavior plays a major role in conveying the mental state of a person, which can be reflected by the movements of facial muscles. Ekman and Fries conducted a cross-cultural study on facial expressions, showing that there are six basic universal facial emotions across human ethnicity and cultures – Anger, Disgust, Fear, Happiness, Sadness, and Surprise [2] (see Figure 1). Subsequently, Contempt has been added to these basic emotions [3]. Given the simplicity of discrete representation, these seven prototypical emotions are the most widely used categorical models for the classification of facial emotions. Though the terms ”expression” and ”emotion” are used interchangeably in many research papers in the literature, the primary difference is facial emotion conveys the mental state of a person, whereas facial expression is the indicator of the emotions being felt, i.e., facial expressions display a wide range of facial modulations but facial emotions are limited.
@@ 1 Introduction
Facial Action Coding System (FACS) is a taxonomy of facial expressions developed by Ekman [6] that defines all observable facial movements for every emotion. It is comprised of 32 action units (AUs), and 14 additional action descriptors (ADs). AUs are described by the fundamental actions of individual muscles or a group of muscles to form a specific movement (see Figure 2), while ADs are unitary movements that account for the head pose, gaze direction, and miscellaneous actions such as jaw thrust, blow, bite, etc. The FACS has been used as a standard for manually annotating facial expressions as it defines a set of rules to express any possible facial expression in terms of specific AUs and also measures the intensity of facial expressions at five discrete levels ($A<B<C<D<E$), where $A$ being the minimum intensity level and $E$ being the maximum intensity level. Continuous models over affect dimensions have also been proposed in order to further enhance the range of facial expressions [7]. Most of the wild datasets for the recognition of seven basic expressions [8, 9] or AUs [10] are provided as images, resulting in a limited number of samples. On the other hand, datasets for regression tasks are provided as videos [11], which can provide a large number of video frames. Therefore, the problem of limited samples is more pronounced in the case of recognizing seven basic expressions or AUs than that of continuous or ordinal regression levels.
Some of the important challenges in developing robust systems for real-world applications are uncontrolled capture conditions (e.g., illumination) and subjective variations of facial behavior that can degrade system performance. Moreover, the subjective differences in visual appearance cause identity bias. Disentangling features linked to identity from features relevant to expression may play a crucial role in developing a robust system for facial behavior analysis. In addition to identity bias, other personal attributes such as age, gender, and ethnicity may further degrade performance. Finally, there is an imbalance and limited availability of relevant image data that elicit expressions or AUs.
@@
Facial expression recognition (FER) has been a topic of significant interest for several decades [5]. Many techniques have been proposed, often to detect the seven universally basic types of emotions [12], but also to estimate the level of valence and arousal [13], or intensity of affective states like fatigue [14], pain [15], and depression [16]. Most of the early approaches for FER rely on deterministic hand-crafted feature descriptors, such as the Local Binary Pattern (LBP) [17] and LBP on three orthogonal planes (LBP-TOP) [18]. Though these descriptors work well for data captured under controlled environments, they fail to handle the wide range of variations in uncontrolled real-time environments, thereby resulting in the decline of system performance. Recent progress in deep learning (DL) architectures and computing capability allowed the development of state-of-the-art FER models to achieve an impressive level of performance, especially by training end-to-end in a supervised fashion on large fully annotated image datasets [12]. However, there is typically a limited amount of relevant labeled data in the public domain to train accurate DL models. Collecting and annotating large-scale datasets for facial expressions or action units is a costly and time-consuming process and requires domain experts. Moreover, manual annotation is a complex process that is subjective in nature, often resulting in ambiguous annotations, especially for intensities and dimensional labels. For instance, in order to achieve minimal competency as a FACS coder, it takes over 100 hours of training, and each minute of video requires approximately one hour to score [19]. Annotating dimensional real-value labels, (e.g., the 2D valence and arousal space) on a continuous scale becomes even more challenging than discrete intensity levels, due to the wide range of emotions.
@@
Given the cost of collecting and annotating data from each operational environment, DL models are commonly pre-trained with a large annotated generic dataset and then fine-tuned on a labeled development dataset having the same task and a similar sample distribution to the operational environment. For real-world applications, system accuracy can decline if there is considerable divergence or shift between the capture conditions in development and operational environments. For accurate FER, DL models should therefore be adapted to specific individuals and capture conditions (sensors, computing device, and environment), using weakly-labeled or unlabeled data from the operational environment. In order to fully leverage the potential of ML/DL models, several specialized weakly-supervised learning (WSL) methods have been proposed for training on images and videos with limited annotations. This paper focuses on state-of-the-art WSL methods for robust facial behavior analysis based on weakly annotated data.
@@
Although several surveys have been published in recent years that focus on state-of-the-art methods for facial behavior analysis, [5, 20, 21, 12, 15, 22], no survey is focused on WSL methods, despite their relevance in real-world scenarios. Most surveys on facial behavior analysis in the literature provide a taxonomy of methods based on the standard algorithmic pipeline and machine learning paradigms [5, 20, 21]. Given the current trends in affective computing, more recent surveys, like Shan et al. [12], focus on DL models for FER, or focus on specific application areas, like the FER for pain intensity estimation [15].
@@
Figure 3: An illustration of WSL scenarios for expression recognition in videos. (a) Baseline supervised learning using frame-level expression labels. (b) Inexact WSL: multiple instances learning (MIL) with sequence-level expression labels. (c) Incomplete WSL: semi-supervised learning (SSL) using partial frame-level expression labels. (d) Inaccurate WSL: supervised learning with noisy expression labels. ${\mathbf{X}}_{1}$, …, ${\mathbf{X}}_{N}$ represents $N$ input video sequences. ${\mathbf{Y}}_{1}$, …, ${\mathbf{Y}}_{N}$ represents the sequence-level labels, while $\widetilde{\mathbf{Y}}_{1}$, …, $\widetilde{\mathbf{Y}}_{N}$ represents the sequence-level model predictions. ${\mathbf{x}}_{11}$, ${\mathbf{x}}_{12}$, …, ${\mathbf{x}}_{1n}$ are the frames of first sequence ${\mathbf{X}}_{1}$, and ${\mathbf{y}}_{11}$, ${\mathbf{y}}_{12}$, …, ${\mathbf{y}}_{1n}$ represents their respective frame-level expression labels. $\widetilde{\mathbf{y}}_{1}$, $\widetilde{\mathbf{y}}_{2}$, …., $\widetilde{\mathbf{y}}_{n}$ denotes frame level predictions. $\overline{\mathbf{y}}_{\mathbf{1}}$, $\overline{\mathbf{y}}_{2}$, …, $\overline{\mathbf{y}}_{n}$ refers to noisy frame-level expression labels. ”?” refers to frames with no labels.
@@
In this paper, we present a comprehensive review of WSL methods that are relevant to various scenarios for both image- and video-based analysis of facial behavior, primarily focusing on classification (using discrete/categorical labels) and regression (using ordinal/dimensional labels) applications. Specifically, we identify various sub-problems of WSL for facial behavior analysis and present the relevant WSL methods along with their advantages and limitations. Given the immense potential of AUs for interpreting expressions, and deriving high-level information, this paper reviews methods for the analysis of both facial expressions and AUs. Cases involving discrete/categorical labels are analyzed as classification or detection problems in expression or AU. Ordinal regression involves the estimation of discrete ordinal or intensity levels of expressions or AUs, whereas dimensional regression involves estimating the wide range of emotions on a continuous real-valued scale in, e.g., the 2-D valence and arousal space. In addition, the performance of WSL methods is compared and interpreted for different scenarios based on standard experimental protocols. Given our detailed critical analysis of WSL methods and their performance, we also discuss the limitations of the state-of-the-art methods, open challenges, and future potential research directions to further advance the development of WSL methods for facial behavior analysis.
@@
The main contributions of this paper are as follows. (1) A taxonomy of WSL scenarios and their relevance in facial behavior analysis, along with the corresponding problem formulations for applications that involve weak discrete-valued labels for classification) and real-valued labels for regression (see Section 2). (2) A comprehensive review of state-of-the-art WSL methods proposed under different scenarios for classification applications (i.e., expression and AU recognition), and for regression applications (i.e., expression and AU intensity estimation), along with their advantages and limitations (see Sections 3 and 4, respectively). (3) An overview of experimental evaluation strategies for evaluation of WSL methods (widely used datasets and protocols), and comparative results with state-of-the-art methods under different WSL scenarios (see Sections 3.4 and 4.3). (4) An analysis of the key challenges with the latest WSL methods and research opportunities associated with the development of robust systems using weakly annotated data in real-world classification and regression applications (see Section 5).
2 Weakly Supervised Learning Scenarios
@@
WSL scenarios can be classified as inexact, incomplete, and inaccurate supervision [23] based on the availability of annotations. This section introduces these three categories, and their relevance in the recognition of facial expressions and AUs, as depicted in Figures 3 and 4, respectively. This paper focuses primarily on four specific applications that are relevant in WSL for facial behavior analysis: expression detection and intensity estimation, and AU detection and intensity estimation.
@@ 2.1 Inexact Supervision
In this category, the labeling is provided for the data samples at a high level (sequence level for expressions or image level for AUs) instead of the dense labeling of the data. The goal is to predict the accurate high-level or low-level labels of unknown test data using the coarsely labeled training data. Multiple Instance Learning (MIL) [24] techniques provide the main approaches to tackle these scenarios. The coarsely labeled data are considered a ”bag”, and the samples within the bag contributing to the coarse annotation are referred to as ”instances”. Several factors influence the performance of MIL algorithms such as bag composition, data distribution, and label ambiguity [25]. The expression detection task seeks to localize and predict the expressions of short video clips or frames using training data with sequence-level labels. This problem can be overcome using MIL algorithms, where sequence-level predictions can be estimated along with frame-level predictions (see Figure 3b). In the context of multiple instance regression (MIR) for expression intensity estimation, the objective is to estimate the intensities of frames or sequences using training data from videos with sequence level labels of expression intensities, where the sequence level label is given by the maximum or average of labels of individual frames of a given sequence [26]. State-of-the-art approaches for detecting facial expressions and expression intensities from sequence-level labels are discussed in Sections 3.1.1 and 4.1.1, respectively.
@@
The task of AU detection in the context of inexact annotations is formulated as the problem of AU estimation from the expression labels of training data without AU labels, where expression labels provide weak supervision for AU labels as shown in Figure 4b. The relevant approaches for the problem of AU detection from the training data of expression labels are elaborated in detail in Section 3.1.2. Similar to MIR for expression intensity estimation, AU intensity estimation can also be formulated in order to estimate the AU intensities of frames (instances) or sequences (bags) using training data of sequence-level AU intensities. State-of-the-art approaches for estimating the intensity of AU are discussed in Section 4.1.2.
2.2 Incomplete Supervision
This category refers to the family of ML algorithms dealing with situations where only a small amount of labeled data is provided, despite the availability of abundant unlabeled data. One of the major solutions to handle this problem is semi-supervised learning (SSL) [27], where data distribution is assumed to occur in clusters.
@@ 2.2 Incomplete Supervision
In the context of incomplete supervision for expression detection in videos (or images), annotations will be provided for a subset of videos (or images). The objective is to predict the labels of all the frames in a test sequence (or test image) using partial labels of videos (or images) of training data. For expression intensity estimation, the intensity levels will be provided only for a subset of videos (or images) as shown in Figure 3c. The relevant approaches for expression detection and intensity estimation are discussed in Section 3.2.1 and Section 4.2.1 respectively. For AU detection, the problem of semi-supervised learning can be formulated in two ways: Missing Labels and Incomplete Labels as shown in Figure 4c. In the first case, each sample of training data is assumed to be provided with multiple labels of Action Units but with missing labels. The task is to train the AU classifier with the training samples of missing labels to predict the complete set of labels for the test sample. On the other hand, for incomplete labels, the entire label set of multiple AUs is provided to the training data but only for a subset of images in the training data, where the rest of the training samples do not have AU labels (but may have expression labels). The problem of incomplete annotation can be extended further to AU intensity estimation, where AU intensity levels are provided only for the keyframes within a video sequence of the training data or only for a subset of the dataset of images. The relevant approaches for AU detection and AU intensity estimation are discussed in Section 3.2.2 and Section 4.2.2 respectively.
2.3 Inaccurate Supervision

### v4

@@ Abstract
Given the recent advancements in deep learning (DL), and computing capabilities, there has been significant progress in the development of systems analyzing facial behavior, evolving from images under controlled laboratory conditions, to more challenging real-world scenarios, including video-based recognition. However, DL models typically require training with large-scale datasets to provide a high level of performance. The collection and annotation of such datasets is a costly undertaking that relies on domain experts. Moreover, the annotation process is highly vulnerable to the ambiguity of expressions or Action Units (AUs) due to the bias induced by experts. This paper provides a comprehensive review of state-of-the-art weakly supervised learning (WSL) methods for facial behavior analysis based on images and videos. First, we introduce a taxonomy of relevant WSL scenarios, and then methods proposed for classification applications (using weak categorical/discrete labels), and for regression applications (using weak ordinal/dimensional labels). For both types of recognition applications (classification and regression), we provide a systematic review of state-of-the-art machine learning (ML) models for expression and AU recognition, and intensity estimation in each WSL scenario, discussing their strengths and limitations. An overview of widely used experimental methodologies (public datasets and protocols) is also provided for the performance evaluation of these state-of-the-art models. Finally, our critical analysis of the methods, along with their experimental results, provides insight into the key challenges of WSL in different scenarios, motivating future research to leverage weakly labeled data for real-world facial behavior analysis problems. Our review indicates that WSL methods may provide a cost-effective approach to train robust ML/DL models for facial behavior analysis in real-world scenarios. A website including constantly-updated survey is provided at https://github.com/praveena2j/awesome-Weakly-Supervised-Facial-Behavior-Analysis.
@@
Weakly Supervised Learning, Facial Expression Recognition, Intensity Estimation, Action Unit Detection, Deep Learning.
I Introduction
Facial Behavior Analysis is an emerging area of interest in computer vision and affective computing, with immense potential for many applications in human-computer interaction, sociable robots, autonomous-driving, etc. It was shown that only one-third of human communication is conveyed through verbal components and two-thirds of communication occurs through non-verbal components [1]. Although emotions can be conveyed through multiple modalities, facial behavior plays a major role in conveying the mental state of a person, which can be reflected by the movements of facial muscles. Ekman and Fries conducted a cross-cultural study on facial expressions, showing that there are six basic universal facial emotions – Anger, Disgust, Fear, Happiness, Sadness, and Surprise [2] (see Figure 1). Subsequently, Contempt has been added to these basic emotions [3]. Given the simplicity of discrete representation, these seven prototypical emotions are the most widely used categorical models for the classification of facial emotions. Though the terms ”expression” and ”emotion” are used interchangeably in the literature, the primary difference is facial emotion conveys the mental state of a person, whereas facial expression is the indicator of the emotions being felt, i.e., facial expressions display a wide range of facial modulations but facial emotions are limited.
@@ I Introduction
Facial Action Coding System (FACS) is a taxonomy of facial expressions developed by Ekman [6] that defines all observable facial movements for every emotion. It is comprised of 32 action units (AUs), and 14 additional action descriptors (ADs). AUs are described by the fundamental actions of individual muscles or a group of muscles to form a specific movement (see Figure 2), while ADs are unitary movements that account for the head pose, gaze direction, and miscellaneous actions such as jaw thrust, blow, bite, etc. The FACS has been used as a standard for manually annotating facial expressions as it defines a set of rules to express any possible facial expression in terms of specific AUs and also measure the intensity of facial expressions at five discrete levels ($A<B<C<D<E$). Continuous models over affect dimensions have also been proposed to further enhance the range of facial expressions [7]. Most of the wild datasets for the recognition of seven basic expressions [8, 9] or AUs [10] are provided as images, resulting in a limited number of samples. On the other hand, datasets for regression tasks are provided as videos [11], which can provide a large number of video frames. Therefore, the problem of limited samples is more pronounced in the case of recognizing seven basic expressions or AUs than that of continuous or ordinal regression levels.
@@
Several techniques have been proposed for facial expression recognition (FER), often to detect the seven universal basic emotions [12], but also to estimate the level of valence and arousal [13], or intensity of affective states like fatigue [14], pain [15], and depression [16]. Most of the early approaches for FER rely on deterministic hand-crafted feature descriptors, such as the Local Binary Pattern (LBP) [17] and LBP on three orthogonal planes (LBP-TOP) [18]. Though these descriptors work well for data captured under controlled environments, they fail to handle the wide range of variations in uncontrolled real-time environments due to limited representation capability. Recent progress in deep learning (DL) architectures and computing capability allowed the development of state-of-the-art FER models to achieve an impressive level of performance, especially by training end-to-end in a supervised fashion on large-scale fully annotated datasets [12]. However, there is typically a limited amount of relevant labeled data in the public domain, which becomes a bottleneck to accurately train data-hungry DL models. Collecting and annotating large-scale datasets for facial expressions or action units is a costly and time-consuming process. Moreover, manual annotation is a complex process that requires domain experts, which may be subjective in nature, often resulting in ambiguous annotations, especially for intensities and dimensional labels. For instance, in order to achieve minimal competency as a FACS coder, it takes over 100 hours of training, and each minute of video requires approximately one hour to annotate [19]. Annotating dimensional real-value labels, (e.g., the 2D valence and arousal space) on a continuous scale becomes even more challenging than discrete intensity levels, due to the wide range of emotions.
@@
Given the cost of collecting and annotating data from each operational environment, DL models are commonly pre-trained with a large annotated generic dataset and then fine-tuned on the limited labeled dataset of the operational environment. For real-world applications, system accuracy can decline if there is considerable divergence or shift between the capture conditions in development and operational environments. For accurate FER, DL models should therefore be adapted to specific individuals and capture conditions (sensors, computing device, and environment), using weakly-labeled or unlabeled data from the operational environment. In order to fully leverage the potential of ML/DL models, several specialized weakly-supervised learning (WSL) methods have been proposed for training on images and videos with limited annotations. This paper focuses on state-of-the-art WSL methods for robust facial behavior analysis based on weakly annotated data.
@@
Although several surveys have been published that focus on state-of-the-art methods for facial behavior analysis, [5, 20, 21, 12, 15, 22], no survey is focused on WSL methods, despite their relevance in real-world scenarios. Most surveys on facial behavior analysis provide a taxonomy of methods based on the standard algorithmic pipeline and ML paradigms [5, 20, 21]. Given the current trends in affective computing, more recent surveys focused on DL models for FER [12], or specific application areas like pain intensity estimation [15].
@@
Figure 3: An illustration of WSL scenarios for expression recognition in videos. (a) Baseline supervised learning using frame-level expression labels. (b) Inexact WSL: MIL with sequence-level expression labels. (c) Incomplete WSL: semi-supervised learning (SSL) using partial frame-level expression labels. (d) Inaccurate WSL: supervised learning with noisy expression labels. ${\mathbf{X}}_{1}$, …, ${\mathbf{X}}_{N}$ represents $N$ input video sequences. ${\mathbf{Y}}_{1}$, …, ${\mathbf{Y}}_{N}$ represents the sequence-level labels, while $\widetilde{\mathbf{Y}}_{1}$, …, $\widetilde{\mathbf{Y}}_{N}$ represents the sequence-level model predictions. ${\mathbf{x}}_{11}$, ${\mathbf{x}}_{12}$, …, ${\mathbf{x}}_{1n}$ are the frames of first sequence ${\mathbf{X}}_{1}$, and ${\mathbf{y}}_{11}$, ${\mathbf{y}}_{12}$, …, ${\mathbf{y}}_{1n}$ represents their respective frame-level expression labels. $\widetilde{\mathbf{y}}_{1}$, $\widetilde{\mathbf{y}}_{2}$, …., $\widetilde{\mathbf{y}}_{n}$ denotes frame level predictions. $\overline{\mathbf{y}}_{\mathbf{1}}$, $\overline{\mathbf{y}}_{2}$, …, $\overline{\mathbf{y}}_{n}$ refers to noisy frame-level expression labels. ”?” refers to frames with no labels.
@@
In this paper, we present a comprehensive review of WSL methods that are relevant to various scenarios for both image- and video-based analysis of facial behavior, primarily focusing on classification (using discrete/categorical labels) and regression (using ordinal/dimensional labels) applications. Specifically, we identify various sub-problems of WSL for facial behavior analysis and present the relevant WSL methods along with their advantages and limitations. Given the potential of AUs for interpreting expressions, and deriving high-level information, this paper reviews methods analyzing both facial expressions and AUs. Cases involving discrete/categorical labels are analyzed as classification or detection problems in expression or AU. Ordinal regression involves the estimation of discrete ordinal or intensity levels of expressions or AUs, whereas dimensional regression involves estimating the wide range of emotions on a continuous real-valued scale such as 2-D valence and arousal space. The performance of WSL methods is further compared and interpreted for different scenarios based on standard experimental protocols. Given our detailed critical analysis of WSL methods and their performance, we also discuss the limitations of the state-of-the-art methods, open challenges, and future research directions to further advance the development of WSL methods for facial behavior analysis.
@@
The main contributions of this paper can be summarized as: (1) A taxonomy of WSL scenarios and their relevance in facial behavior analysis, along with the corresponding problem formulations for applications related to weak discrete-valued labels for classification and weak real-valued labels for regression (see Section II). (2) A comprehensive review of state-of-the-art WSL methods proposed under different scenarios for classification (i.e., expression and AU recognition), and for regression (i.e., expression and AU intensity estimation), along with the critical analysis of their strengths and limitations (see Sections III and IV, respectively). (3) An overview of experimental protocols for evaluating the WSL methods on widely used datasets, and the comparative results of these methods under different WSL scenarios (see Sections III-D and IV-C). (4) An analysis of the open problems, and research opportunities for the development of robust systems using weakly annotated data in real-world classification and regression applications (see Section V).
II Weakly Supervised Learning Scenarios
@@
WSL scenarios can be classified as inexact, incomplete, and inaccurate supervision based on the availability of annotations [23]. This section introduces these three categories, and their relevance in the recognition of facial expressions and AUs, as depicted in Figures 3 and 4, respectively. This paper focuses primarily on four specific applications that are relevant in WSL for facial behavior analysis: expression detection and intensity estimation, and AU detection and intensity estimation.
@@ II-A Inexact Supervision
In this category, the labeling is provided for the data samples at a coarse level (sequence level for expressions or image level for AUs) instead of dense labeling of the data. The goal is to predict the accurate high-level or low-level labels of unknown test data using the coarsely labeled training data. MIL [24] is one of the major approaches to tackle these scenarios. The coarsely labeled data are considered a ”bag”, and the samples within the bag contributing to the coarse annotation are referred to as ”instances”. Several factors influence the performance of MIL algorithms such as bag composition, data distribution, and label ambiguity [25]. The expression detection task seeks to localize and predict the expressions of short video clips or frames using training data with sequence-level labels (see Figure 3b). In the context of multiple instance regression (MIR) for expression intensity estimation, the goal is to estimate the intensities of frames or sequences using training videos with sequence level labels of expression intensities, where the sequence level label is given by the maximum or average of labels of individual frames of a given sequence [26].
@@
The task of AU detection in the context of inexact annotations is formulated as the problem of AU estimation from the expression labels of training data without AU labels, where expression labels provide weak supervision for AU labels as shown in Figure 4b. Similar to MIR for expression intensity estimation, AU intensity estimation can also be formulated in order to estimate the AU intensities of frames (instances) or sequences (bags) using training data of sequence-level AU intensities.
II-B Incomplete Supervision
@@
This category refers to the family of ML algorithms dealing with situations where only a small amount of labeled data is provided, despite the availability of abundant unlabeled data. Two major strategies to handle this problem is semi-supervised learning (SSL) [27] and active learning [28]. In the context of incomplete supervision for expression detection in videos (or images), annotations will be provided for a subset of videos (or images). The objective is to predict the labels of all the frames in a test sequence (or test image) using partial labels of videos (or images) of training data. For expression intensity estimation, the intensity levels will be provided only for a subset of videos (or images) as shown in Figure 3c. For AU detection, the problem can be formulated in two ways: Missing Labels and Incomplete Labels (see Figure 4c). In the first case, each sample of training data is assumed to be provided with multiple labels of Action Units but with missing labels. The task is to train the AU classifier with the training samples of missing labels to predict the complete set of labels for the test sample. On the other hand, for incomplete labels, the entire label set of multiple AUs is provided to the training data but only for a subset of images in the training data, where the rest of the training samples do not have AU labels (but may have expression labels). The problem of incomplete annotation can be extended further to AU intensity estimation, where AU intensity levels are provided only for the keyframes within a video sequence of the training data or only for a subset of the dataset of images.
II-C Inaccurate Supervision